1. A drag racer constantly accelerates at a rate of 39 m/s2 to increase its speed in a short amount of time. If the drag racer started at rest, what distance did it have to travel to reach a velocity of 33 m/s, E?


Include complete solution please



1
Expert's answer
2021-10-15T10:24:39-0400

There is a simple equation that relates the distance, acceleration, and speed for accelerated motion started from rest:


"d=\frac{v^2}{2a}=\frac{33^2}{2·39}=14\text{ m}."